# Resonant-Tunnelling Diode Reservoir Computing System for Image Recognition

### v1

@@ Abstract
As artificial intelligence continues to push into real-time, edge-based and resource-constrained environments, there is an urgent need for novel, hardware-efficient computational models. In this study, we present and validate a neuromorphic computing architecture based on resonant-tunnelling diodes (RTDs), which exhibit the nonlinear characteristics ideal for physical reservoir computing (RC). We theoretically formulate and numerically implement an RTD-based RC system and demonstrate its effectiveness on two image recognition benchmarks: handwritten digit classification and object recognition using the Fruit 360 dataset. Our results show that this circuit-level architecture delivers promising performance while adhering to the principles of next-generation RC—eliminating random connectivity in favour of a deterministic nonlinear transformation of input signals.
@@ I Introduction
The recognition part of such a system is mostly delegated to software running on a digital computer. Typically, the computer must have considerable computational power to store and process information. While using a high-performance workstation is not an issue in a stationary context—such as a research laboratory server room or data centre—this solution is not suitable for autonomous systems, especially airborne drones, where power supply capacity, weight and cost are critical constraints [9, 10].
While it is technically possible to replace the software and digital computer with electronic [11, 12, 5] or optical [13] neural networks, a more viable solution may lie in a hybrid approach. In this method, certain functions of the neural network—such as the nodes and their associated activation functions—are replaced by electronic components or circuits [14]. This neuromorphic approach, which uses brain-inspired principles to design circuits capable of performing computational tasks with significantly greater power efficiency than conventional computers [15, 16, 17], would still rely on a reasonably fast and low-power digital computer for data processing. However, it would be more efficient overall due to the implementation of neurons as electronic circuits [16, 17].
This approach is particularly well suited to physical (hardware) reservoir computing (RC)—a branch of neuromorphic computing that employs the intrinsic dynamics of physical systems to perform efficient computation [18, 19, 20, 21]. In particular, RC systems have proven effective for processing highly nonlinear and chaotic time-series datasets. A key advantage of RC is its low training cost, as it typically requires minimal training data and relies only on linear optimisation of the readout layer [22]. However, conventional RC models depend on randomly initialised weight matrices to define neuron connectivity and input mapping, which often requires careful tuning. To address this challenge, recent research has shown a mathematical equivalence between RC and nonlinear vector auto-regression (NVAR), a formulation that eliminates the need for random matrices, reduces hyperparameter complexity and offers improved interpretability [23, 24]. This development is referred to as next-generation reservoir computing.
@@ II Resonant-Tunnelling Diode as a Computational Reservoir
It is also instructive to comment on the suitability of RTDs as elements of an RC system. While the algorithmic RC architecture and its modifications are well established and understood [22], the debate over the suitability of different physical systems for RC purposes remains ongoing—primarily because virtually any nonlinear dynamical physical system can, in principle, serve as a reservoir, although not all such systems yield efficient computational performance [39, 40, 41].
Apart from their suitability for broader applications in neuromorphic computing [31, 32, 33], RTDs—more specifically, their $I–V$ characteristics—can, in principle, mathematically substitute standard activation functions, potentially yielding improved overall model performance [42, 43]. (The same holds true for the quantum tunnelling effect model, which is relevant to the operation of an RTD [44].) However, in the computational scheme proposed in this work, RTDs substitute the computationally demanding, nonlinear signal processing component of the reservoir computing algorithm with the physical—albeit numerically simulated in this study—response of RTDs (a hardware implementation of an array of RTDs is technologically viable [45, 38]). As a result, the RTD-based RC system promises to be more energy-efficient while delivering robust computational performance, thereby fulfilling one of the main goals of physical reservoir computing [18, 19, 20, 28].
Figure 2: Schematic illustration of the proposed RTD-based RC architecture. An array of RTDs serves as a computational reservoir that facilitates—by virtue of the nonlinear $I–V$ characteristics of the diodes—the transformation of low-dimensional input signals into high-dimensional representations. The reservoir output—the currents $I_{i}$, $i=0\dots n$, produced by each diode—is processed using a linear readout mechanism.
@@ III Reservoir Computing Model
The RTD-based RC system processes grayscale images by converting them into electric voltage signals. To this end, each input image is first transformed into a binary format and then normalised to align with the diode’s operating voltage range, $V\in[-5,5]$. Such an image processing is conducted row-wise: each horizontal row of pixels in the image is encoded into a sequence of voltage pulses applied to a corresponding RTD in the array. Thus, for an image of size $n\times m$ pixels, the reservoir comprises $n$ RTDs, each receiving a pulse train of length $m$ that represents the pixel intensities of its row. For example, with standard MNIST images of $28\times 28$ pixels, the system utilises 28 diodes, each driven by a 28-element voltage pulse sequence. The values $I_{i}$, $i=0\dots n$, of the output current produced by each diode are treated as the dynamic reservoir states [22], which form a high-dimensional embedding of the input and used to train a linear readout (Fig. 2).
@@ IV Results and Discussion
To theoretically evaluate the performance of the RTD-based RC system, the equations of the RTD model and the equations of the RC model presented above were implemented as a software package in Python 3. The resulting software was tested on two benchmark image recognition datasets: the standard MNIST and Fruit-360 [46]. The well-known MNIST dataset contains 60,000 grayscale images of handwritten digits from 0 to 9, each ones sized $28\times 28$ pixels. The Fruit-360 dataset includes 10,000 grayscale images representing 15 fruit classes, resized to $28\times 28$ pixels. Both datasets were split into 70% training and 30% testing subsets.
All numerical experiments were conducted using a 64-bit workstation computer powered by an Apple M3 Max CPU and equipped with 128 GB of unified memory. The input images were shuffled prior to training and testing to ensure unbiased results. The voltage mapping range applied to the RTD array was optimised and set to $[-5,5]$ V.
@@ V Connection to the Next-Generation Reservoir Computing Framework
Furthermore, based on the results presented in Ref. [24], it has been demonstrated that this computational scheme not only circumvents the use of randomly generated neural connection matrices but is also equivalent to—and in some cases more efficient than—the traditional RC algorithm. This approach is referred to as next-generation reservoir computing [23, 50].

### v2

@@ Abstract
As artificial intelligence continues to push into real-time, edge-based and resource-constrained environments, there is an urgent need for novel, hardware-efficient computational models. In this study, we present and validate a neuromorphic computing architecture based on resonant-tunnelling diodes (RTDs), which exhibit the nonlinear characteristics ideal for physical reservoir computing (RC). We theoretically formulate and numerically implement an RTD-based RC system and demonstrate its effectiveness on two image recognition benchmarks: handwritten digit classification and object recognition using the Fruit 360 dataset. Our results show that this circuit-level architecture delivers promising performance while adhering to the principles of next-generation RC, eliminating random connectivity in favour of a deterministic nonlinear transformation of input signals.
@@ I Introduction
The recognition part of such a system is mostly delegated to software running on a digital computer. Typically, the computer must have considerable computational power to store and process information. While using a high-performance workstation is not an issue in a stationary context, such as a research laboratory server room or data centre, this solution is not suitable for autonomous systems, especially airborne drones, where power supply capacity, weight and cost are critical constraints [9, 10].
While it is technically possible to replace the software and digital computer with electronic [11, 12, 5] or optical [13] neural networks, a more viable solution may lie in a hybrid approach. In this method, certain functions of the neural network, such as the nodes and their associated activation functions, are replaced by electronic components or circuits [14]. This neuromorphic approach, which uses brain-inspired principles to design circuits capable of performing computational tasks with significantly greater power efficiency than conventional computers [15, 16, 17], would still rely on a reasonably fast and low-power digital computer for data processing. However, it would be more efficient overall due to the implementation of neurons as electronic circuits [16, 17].
This approach is particularly well suited to physical (hardware) reservoir computing (RC), a branch of neuromorphic computing that employs the intrinsic dynamics of physical systems to perform efficient computation [18, 19, 20, 21]. In particular, RC systems have proven effective for processing highly nonlinear and chaotic time-series datasets. A key advantage of RC is its low training cost, as it typically requires minimal training data and relies only on linear optimisation of the readout layer [22]. However, conventional RC models depend on randomly initialised weight matrices to define neuron connectivity and input mapping, which often requires careful tuning. To address this challenge, recent research has shown a mathematical equivalence between RC and nonlinear vector auto-regression (NVAR), a formulation that eliminates the need for random matrices, reduces hyperparameter complexity and offers improved interpretability [23, 24]. This development is referred to as next-generation reservoir computing.
@@ II Resonant-Tunnelling Diode as a Computational Reservoir
It is also instructive to comment on the suitability of RTDs as elements of an RC system. While the algorithmic RC architecture and its modifications are well established and understood [22], the debate over the suitability of different physical systems for RC purposes remains ongoing, primarily because virtually any nonlinear dynamical physical system can, in principle, serve as a reservoir, although not all such systems yield efficient computational performance [39, 40, 41].
Apart from their suitability for broader applications in neuromorphic computing [31, 32, 33], RTDs (more specifically, their $I–V$ characteristics) can, in principle, mathematically substitute standard activation functions, potentially yielding improved overall model performance [42, 43]. (The same holds true for the quantum tunnelling effect model, which is relevant to the operation of an RTD [44].) However, in the computational scheme proposed in this work, RTDs substitute the computationally demanding, nonlinear signal processing component of the reservoir computing algorithm with the physical—albeit numerically simulated in this study—response of RTDs (a hardware implementation of an array of RTDs is technologically viable [45, 38]). As a result, the RTD-based RC system promises to be more energy-efficient while delivering robust computational performance, thereby fulfilling one of the main goals of physical reservoir computing [18, 19, 20, 28].
Figure 2: Schematic illustration of the proposed RTD-based RC architecture. An array of RTDs serves as a computational reservoir that facilitates, by virtue of the nonlinear $I–V$ characteristics of the diodes, the transformation of low-dimensional input signals into high-dimensional representations. The reservoir output (the currents $I_{i}$, $i=0\dots n$, produced by each diode) is processed using a linear readout mechanism.
@@ III Reservoir Computing Model
The RTD-based RC system processes grayscale images by converting them into electric voltage signals. To this end, each input image is first transformed into a binary format and then normalised to align with the diode’s operating voltage range, $V\in[-5,5]$. Such an image processing is conducted row-wise: each horizontal row of pixels in the image is encoded into a sequence of voltage pulses applied to a corresponding RTD in the array. Thus, for an image of size $n\times m$ pixels, the reservoir comprises $n$ RTDs, each receiving a pulse train of length $m$ that represents the pixel intensities of its row. For example, with standard MNIST images of $\times 28$ 28 pixels, the system utilises 28 diodes, each driven by a 28-element voltage pulse sequence. The values $I_{i}$, $i=0\dots n$, of the output current produced by each diode are treated as the dynamic reservoir states [22], which form a high-dimensional embedding of the input and used to train a linear readout (Fig. 2).
@@ IV Results and Discussion
To theoretically evaluate the performance of the RTD-based RC system, the equations of the RTD model and the equations of the RC model presented above were implemented as a software package in Python 3. The resulting software was tested on two benchmark image recognition datasets: the standard MNIST and Fruit-360 [46]. The well-known MNIST dataset contains 60,000 grayscale images of handwritten digits from 0 to 9, each ones sized $\times 28$ 28 pixels. The Fruit-360 dataset includes 10,000 grayscale images representing 15 fruit classes, resized to $\times 28$ 28 pixels. Both datasets were split into 70% training and 30% testing subsets.
All numerical experiments were conducted using a 64-bit workstation computer powered by an Apple M3 Max CPU and equipped with 128 GB of unified memory. The input images were shuffled prior to training and testing to ensure unbiased results. The voltage mapping range applied to the RTD array was optimised and set to [5,5][-5,5][ - 5 , 5 ] V.
@@ V Connection to the Next-Generation Reservoir Computing Framework
Furthermore, based on the results presented in Ref. [24], it has been demonstrated that this computational scheme not only circumvents the use of randomly generated neural connection matrices but is also equivalent to, and in some cases more efficient than, the traditional RC algorithm. This approach is referred to as next-generation reservoir computing [23, 50].